OpenStudy (anonymous):

Directions: Solve the problems below. Make sure to show your work when applicable. Simplify the expression. 22+(3^2-4^2) Simplify the expression. √625 Evaluate the expression for a = –1 and b = 5. 5a – 7 Evaluate the expression for a = –1 and b = 5. (ab)2 Simplify the expression. 9xy2 – 11xy2 Simplify the expression. -3(b – 7) Solve the equation. Check your answer. m/20 = 5/4 Solve the equation. Check your answer. 4y + 5 = – 31 Solve the equation. Check your answer. 3x – 7 = 5x + 19

jagr2713 (jagr2713):

There isnt a b for the 3seond question

jagr2713 (jagr2713):

Simplify the expression. -3(b – 7) Distribute and get 21-3b

OpenStudy (anonymous):

I know but that's the exact question...

jagr2713 (jagr2713):

Solve the equation. Check your answer. m/20 = 5/4 Cross multiply 4*20 and 5*m 80/5m divide and get m=16

jagr2713 (jagr2713):

Solve the equation. Check your answer. 4y + 5 = – 31 subtract 5 on both sides 4y=-36 then divide by 4 y=-9

jagr2713 (jagr2713):

Solve the equation. Check your answer. 3x – 7 = 5x + 19 start by finding like terms 3x-5=-2 -2x-7=19 +7 +7 -2x=26 divide x=-13

jagr2713 (jagr2713):

Evaluate the expression for a = –1 and b = 5. (ab)2 put the numbers into the equation (ab)^2, (-1*5)^2, (-5)^2, (-5)(-5), 25.

jagr2713 (jagr2713):

Simplify the expression. 9xy2 – 11xy2 9xy² - 11xy² = -2xy²

jagr2713 (jagr2713):

Simplify the expression. 22+(3^2-4^2) start within the parentheses 22+( 9-16) 22+(-7) add and get 15

jagr2713 (jagr2713):

Simplify the expression. √625 all you do is put this in a calc and thats it so you can do this

jagr2713 (jagr2713):

I think thats all :D